BAT Industries is a manufacturer of orange juice in Jamaica. In 2022 it sold a bottle of orange juice to wholesalers for $100 and its variable costs is $40 and fixed costs is $100,000 monthly:

 

 

Using the information provided. 

 

i. How many bottles of orange juice must the company sell in order to break even and what is the breakeven revenue? (4 marks)

 

 

 

ii. If the company sold 1,500,000 bottles of orange juice for the year 2022, what was its profit? (4 marks)

 

 

 

iii. Due to inflation, in 2023 fixed costs are expected to rise to $175,000 monthly. What will be the NEW break-even quantity and revenue? (4 marks)

 

 

 

iv. If the company will sell the same number of orange juice as it did in 2022 and it wants to make the same profit as in 20222, what price should they sell each bottle of orange Juice for in 2023? (5 marks)

1
Expert's answer
2022-03-18T12:41:14-0400

i

n - number of sold bottles.

Total production cost is 40*n + 100000

Total revenue is 100*n

The breakeven quantity defined by 40*n+100000 = 100*n

n = 100000/60 = 1667

Total revenue is 100*n = 166700


ii

The profit is 100*n - 40*n - 100000 = 60*n - 100000.

For n = 1,500,000 the profit is 60*1,500,000 - 100,000 = 90,000,000 - 100,000 = 89,900,000


iii

New total production cost is 40*n + 175,000

Total revenue stiдl is 100*n

The breakeven quantity defined by 40*n+175,000 = 100*n

n = 175,000/60 = 2,917

Total revenue is 100*n = 291,700


iv

The new price should compensate the growth of fixed cost: 175,000 - 100,000 = 75,000.

So the price of a single bottle should increase by 75,000 / 1,500,000 = 0.05,

and the new price should be 100.05